OpenStudy (anonymous):

Solve x/2- 3 = 7. a.5 b.10 c.20 d.40

OpenStudy (anonymous):

@kaiserTheSlayer @rahammi

OpenStudy (anonymous):

@Rahammii

OpenStudy (anonymous):

okay so you first need to subtract the two. 2-3 is what?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

okay so what times seven equals seven?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

exactly. division is the oppisate of multiplication. 1*7=7. so now what do you think x equals?

OpenStudy (anonymous):

7-2=5

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

okay wait.

OpenStudy (anonymous):

what is it then

OpenStudy (anonymous):

the answers given do not make sense with the problem. the answer for x should be seven because 7/1=7 therefore x should be 7 but there isn't that choice so that is kind of confusing.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

is the problem \[\frac{x}{2}- 3 = 7 \]?

OpenStudy (anonymous):

x=20

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (anonymous):

is it 5

OpenStudy (anonymous):

Let's solve your equation step-by-step. x/2−3=7 Step 1: Simplify both sides of the equation. x/2−3=7 Let's simplify step-by-step. x/2−3=7 1/2x+−3=7 1/2x−3=7 Step 2: Add 3 to both sides. 1/2x−3+3=7+3 1/2x=10 Step 3: Multiply both sides by 2. 2*(1/2x)=(2)*(10) x=20